OpenStudy (anonymous):

What is the arc measure of an arc with length 4.189 cm and radius equal to 3 cm. A. 480° B. 53° C. 82° D. 80°

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (lovelycharm):

L = r* theta where theta = arc measure theta = L / r = 4.189 / 3 radians multiply this by 180/pi gives the arc measure in degrees this comes to 80 degrres D

OpenStudy (anonymous):

\[\frac{ \angle measure }{ 360 } = \frac{ arc length }{ circumference }\]and your circumference is: (2)(radius)(pi)

OpenStudy (anonymous):

Rearranging and simplifying:\[\angle measure = \frac{ arclength \times 360 }{ 2 \times radius \times pi }\]

OpenStudy (anonymous):

You have arclength and radius, and you can use 3.14 for pi. At this point, it's just simple arithmetic.

OpenStudy (anonymous):

All good now, @katlin95 ?

OpenStudy (anonymous):

I got 80 as an answer:)

OpenStudy (anonymous):

You are right! Good job!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @katlin95